----------------------------------------------------------------------------- Mauna Loa Solar Observatory Observer's Log ----------------------------------------------------------------------------- Wed Oct 21 15:40:00 GMT 2026 Year:26 Doy:294 Observer: lisapg Weather COMMENT BY lisapg: Wed Oct 21 16:16:11 GMT 2026 MLO Current Weather Timestamp (UTC): 2026-10-21 16:11:00 Temperature : 34.64 deg F Dew Point : 2.26 deg F Pressure : 679.32 mbar Wind Speed : 6.93 mph Wind Direction : 174.3 deg Skies: Sky looks clear ___end___ GENERAL COMMENT BY lisapg: Wed Oct 21 16:35:32 GMT 2026 Construction crew on site ___end___ GENERAL COMMENT BY lisapg: Wed Oct 21 17:07:24 GMT 2026 Dome open ___end___ ** UCOMP PM lisapg ** : Wed Oct 21 17:07:29 GMT 2026 Blew off UCoMP O1 ___end___ ** KCOR PM lisapg ** : Wed Oct 21 17:07:34 GMT 2026 Blew off Kcor O1 ___end___ KCOR COMMENT BY lisapg: Wed Oct 21 17:19:14 GMT 2026 Kcor above dome ___end___ Wed Oct 21 17:21:53 GMT 2026 Kcor Focus/alignment program exited Wed Oct 21 17:22:56 GMT 2026 KCOR Start Synoptic Patrol GENERAL COMMENT BY berkey: Wed Oct 21 18:14:17 GMT 2026 Watchdog failed twice to send images back to boulder. Restarted it and am running catchup to send what is missing to boulder. ___end___ Wed Oct 21 18:44:06 GMT 2026 KCOR End Patrol GENERAL COMMENT BY berkey: Wed Oct 21 18:46:24 GMT 2026 Dome closed wind shifted and was seeing some dust kick up. Will start up again soon, hopefully ___end___ Wed Oct 21 19:12:17 GMT 2026 Kcor Focus/alignment program exited Wed Oct 21 19:12:59 GMT 2026 KCOR Start Synoptic Patrol Wed Oct 21 19:22:14 GMT 2026 KCOR End Patrol Wed Oct 21 19:22:15 GMT 2026 KCOR Start Calibration script: c:\kcor\mlso-calibration22deg-20171025.ini Wed Oct 21 19:37:29 GMT 2026 KCOR End Calibration Script Wed Oct 21 19:37:46 GMT 2026 KCOR Start Synoptic Patrol Wed Oct 21 19:37:47 GMT 2026 KCOR Start Synoptic Patrol GENERAL COMMENT BY berkey: Wed Oct 21 20:46:24 GMT 2026 Construction crew has let ___end___ Wed Oct 21 21:36:05 GMT 2026 KCOR End Patrol GENERAL COMMENT BY berkey: Wed Oct 21 22:00:17 GMT 2026 KCor was stopped for some dome camera work; clouds have rolled in now, waiting for them to dissipate before starting up again. ___end___ Wed Oct 21 22:05:39 GMT 2026 Kcor Focus/alignment program exited Wed Oct 21 22:06:49 GMT 2026 KCOR Start Synoptic Patrol Wed Oct 21 22:08:02 GMT 2026 KCOR End Patrol GENERAL COMMENT BY berkey: Wed Oct 21 22:24:21 GMT 2026 entry was edited Dome closed for dark clouds ___end___ UCOMP COMMENT BY berkey: Wed Oct 21 23:45:17 GMT 2026 Testing a different vga splitter for UComp computer as we were having a display resolution with the one we normally use. ___end___ KCOR COMMENT BY KCOR: Wed Oct 21 23:46:31 GMT 2026 KCOR file transfer report KCor created 944 .bin files 944 .fts files Remote site is complete ? no rsync required. ___end___ GENERAL COMMENT BY berkey: Wed Oct 21 23:55:54 GMT 2026 A nice day to take data; unfortunately clouds moved in for the rest of the day. ___end___ Logger COMMENT BY berkey: Wed Oct 21 23:56:01 GMT 2026 Log closed ONSITE STAFF: berkey, lisapg ___end___